Vector A has magnitude 6.00 m and vector has magnitude 3.00 m. The vector product between these two vectors has magnitude 12,0 square meters. What are the two possible values for the scalar product of these two vectors?


1
Expert's answer
2021-03-04T11:53:46-0500

According to the definition, the vector product gives us the magnitude of the vector and its direction:


A×B=ABsinθ.A\times B=||A|||B||sin\theta.

Let's find the angle θ\theta between the two vectors:


θ=sin1(A×BAB)=sin1(12 m26 m3 m)=41.8.\theta=sin^{-1}(\dfrac{A\times B}{||A|||B||})=sin^{-1}(\dfrac{12\ m^2}{6\ m\cdot3\ m})=41.8^{\circ}.


If our vectors located in the second quadrant, for example, the angle between them will be:


θ=18041.8=138.2.\theta=180^{\circ}-41.8^{\circ}=138.2^{\circ}.


Then, we can find the two possible values for the scalar product of these two vectors:


AB=6 m3 mcos41.8=13.42 m2,A\cdot B=||6\ m|||3\ m||cos41.8^{\circ}=13.42\ m^2,AB=6 m3 mcos138.2=13.42 m2.A\cdot B=||6\ m|||3\ m||cos138.2^{\circ}=-13.42\ m^2.